OpenStudy (zenmo):

Eliminate the parameter to find a Cartesian equation of the curve for −7 ≤ y ≤ −3

OpenStudy (zenmo):

x = 1 − t^2, y = t − 5, −2 ≤ t ≤ 2

OpenStudy (triciaal):